OpenStudy (anonymous):

A photon of light has a wavelength of about 800nm (1nm= 1*10^-9 m). What is the energy of the photon of light?

OpenStudy (shamim):

Energy of photon=h*f=h*c/lamda=?

OpenStudy (shamim):

Given lamda=800nm=800*10^-9m h=planks constant

OpenStudy (shamim):

Feel free to ask for better understanding

OpenStudy (anonymous):

The only question I have is what are the values that I need to plug into the equation?

OpenStudy (anonymous):

You plug it in the equation where you would usually put an equation @joem820

OpenStudy (anonymous):

What I mean is given the equation E=h*c/lamada , how do i determine the values to put into the equation?

OpenStudy (shamim):

I already told the value of lamda

OpenStudy (shamim):

h=6.63*10^-34j.s (Check this value of h, may b little different value is correct one)

OpenStudy (shamim):

c=3*10^-8m/sec=light velocity

OpenStudy (shamim):

Response plz for better help